OpenStudy (anonymous):

solve. e^-2x=0.3 help please!

jimthompson5910 (jim_thompson5910):

e^(-2x) = 0.3 ln( e^(-2x) ) = ln(0.3) -2x*ln( e ) = ln(0.3) -2x*1 = ln(0.3) -2x = ln(0.3) Keep going to solve for x

OpenStudy (anonymous):

i got X=.60198

OpenStudy (anonymous):

well rounded .6020 that look about right?

jimthompson5910 (jim_thompson5910):

I'm getting that too

OpenStudy (anonymous):

ok awsome thank you

jimthompson5910 (jim_thompson5910):

if you wanted the approximate answer, then it would be that if you want the exact answer, then it's just \[\Large x = -\frac{\ln(0.3)}{2}\]

OpenStudy (anonymous):

ok thank you the above one was fine tho i appreciate your help

jimthompson5910 (jim_thompson5910):

np